FMARS-15 Science Report 26-07-2023

 In Science Report

Submitted by Science Officer Olivia Drayson

  1. EVA:  One EVA took place today, patterned ground was reached and geological and biological samples were taken and brought back to the hab for analysis. 
  1. Algal Investigation: The samples collected during EVA2 including biofilm, algae, moss, water and snow were moved from short-term to long-term storage and were inspected visually with the naked eye and underneath a microscope. The algae was determined to be healthy in the long-term storage location and the home for several worm-like larvae. The species of larvae is yet to be determined. The algal health will be regularly monitored and more study under the microscope will be undertaken at higher magnifications.
  1. Microplastics Study: A stock of Nile Red solution was made using acetone (1mg/mL) and stored in the hab laboratory. The fluorescent microscope was set up and tested and a positive control vial containing shavings of plastic was made up to test the nile red concentration and fluorescent microscope settings.
  1. GPS Route Measurements:  EVA3 took a Garmin GPS device with them to map their route.
  1. Human Factors Study: Heart rate variability was tracked for all crew members using wearable Garmin watches as a marker for stress. Two meditation sessions were conducted per day on 24th, 25th and 26th July with the whole crew. The crew also completed the weekly questionnaire and facial emotion recognition test.
  1. Crew comfort, health and safety: Measurements of indoor air quality were taken due to concerns over rising CO2 levels on the upper floor. Measurements were taken outside, on both levels and in the bathroom. It was suspected that using an old gasoline tank as a urine collection tank was causing the high Benzene levels and the Benzene concentration in the bathroom supported this hypothesis. The ventilation in the bathroom was kept on most of the day and the excess gasoline has been removed. However, the CO2 levels rising each night are most likely due to the crew members respiration into a warm, humid, upper level with improper ventilation. A fan was installed on the attic to help circulate the air within the hab and the doors have been left open most of the day. 
  1. Data loggers: The data loggers were deployed on 23rd July and are being checked visually from the hab daily. The snow run-off at the upstream station has already started to dry off and the data will be checked towards the end of the mission. A solar logger was installed by Andrew on the day of arrival 15th July and has been regularly inspected and charged every 36-48 hours and then repositioned outside of the hab. Passive area radiation dosimeters were temporarily affixed in three locations on 20th July, one on a guy wire outside the hab, one on door to the airlock within the hab and another interior one on the same wall but on the upper level. 

Start typing and press Enter to search